# SN 2021lwz: Another Exotic Luminous and Fast Evolving Optical Type Ic Broad-Lined Supernova ?

### v1

@@ 5.8 Comparisons with fast evolving luminous Type Ic-BL SN 2018gep
All in all, all these comparisons suggest that SN 2021lwz is not a fast-rising hydrogen-poor SLSN (He-rich or not), and neither an LFBOT, or a typical H-poor SESN, but rather pertain to a distinct class of objects.
@@ 5.9 Prospects and final remarks
An interesting test of the magnetar scenario for luminous, fast evolving objects is non-thermal emission. These objects could emit luminous X-ray (Kotera et al. 2013; Metzger et al. 2014; Murase et al. 2015) and radio emission (Murase et al. 2015; Omand et al. 2018; Eftekhari et al. 2021; Murase et al. 2021; Omand et al. 2025a), similar to PTF 10hgi (Eftekhari et al. 2019; Law et al. 2019; Mondal et al. 2020) or SN 2012au (Stroh et al. 2021) after a few years, but on timescales of a few weeks to a year. There could also be bright infrared emission on short timescales due to dust heated by the PWN (Omand et al. 2019), although this emission would likely fade quickly as the dust cools. The dominant non-thermal component can be predicted, in principle, from the nebular spectrum of the supernova, since line emission is sensitive to the underlying PWN spectrum (Omand & Jerkstrand 2023); however, these models are not currently capable of reproducing nebular spectra on timescales of less than a year.
The multi-dimensional structure of these magnetar-driven supernovae could show strong clumping due to Rayleigh-Taylor instabilities between the nebula and ejecta (Blondin & Chevalier 2017; Suzuki & Maeda 2017). When , such as in SN 2021lwz, the shock can break through the ejecta (Blondin & Chevalier 2017), leading to a decrease in the ejecta acceleration efficiency. This process, known as blowout, has been inferred to happen in the Crab Nebula (Omand et al. 2025b). Clumpy ejecta can cause polarization if the clump size is large enough (Tanaka et al. 2017), but it is currently unknown how to determine the clump size distribution from the inferred engine properties of the supernova.
It has been proposed by Zou & Cheng (2018) that Type I SLSNe and Type Ic-BL could be produced by jets. If so, they would both show strong polarization from early times. We point out that for SN 2021lwz our non-detections contribute to rule out this scenario. Additionally, if the energy was injected into the ejecta early, most of it would accelerate the ejecta instead of contributing to the luminosity (Suzuki & Maeda 2021).
@@ 6 Conclusions
An extensive analysis and modeling of the photometry, spectroscopy and polarimetry data of the SN, followed by comparisons with SNe of well know types and with rarer transients, consistently leads to the following results:
@@
Spectroscopically, SN 2021lwz looks like a Type Ic-like SN before maximum light. After maximum light it looks similar to a Type Ic-BL.
@@
The light curve and multi-band light curve modeling of SN 2021lwz suggest it shares similarities with iPTF 16asu, and with SN 2018gep. The three sources show continuum blue dominated spectra before maximum light, even though SN 2021lwz shows spectral features pointing to a Type Ic-like SNe, while iPTF 16asu and SN 2018gep don’t. A possible explanation is that iPTF 16asu and SN 2018gep are more energetic explosions quickly rising to higher photosphere temperatures than is SN 2021lwz.
@@
Acknowledgements
The authors would like to thank Sheng Yang for helpful discussions, and Rosa Clavero and David Morate for their kind support on the NOT. We thank George Khorenzhev (Space Research Institute, Russian Academy of Sciences) for contributing to some observations.
F.P. acknowledges support from the Spanish Ministerio de Ciencia, Innovación y Universidades (MICINN) under grant numbers PID2022-141915NB-C21. R.K.T. has been supported by the NKFIH/OTKA FK-134432 grant of the National Research, Development and Innovation Office of Hungary (NKFIH). C. M. B. O. acknowledges support from the Royal Society (grant Nos. DHF-R1-221175 and DHF-ERE-221005). N. Sarin acknowledges support from the Kavli foundation.
T.-W.C. acknowledges the Yushan Fellow Program by the Ministry of Education, Taiwan for the financial support (MOE-111-YSFMS-0008-001-P1). T.K. acknowledges support from the Research Council of Finland project 360274. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute of Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, IN2P3, University of Warwick, Ruhr University Bochum, Cornell University, and Northwestern University. Operations are conducted by COO, IPAC, and UW. Zwicky Transient Facility access was supported by Northwestern University and the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA). The ZTF forced-photometry service was funded under the Heising-Simons Foundation grant No. 12540303 (PI: Graham).
SED Machine is based upon work supported by the National Science Foundation under Grant No. 1106171.
The Gordon and Betty Moore Foundation, through both the Data-Driven Investigator Program and a dedicated grant, provided critical funding for SkyPortal.
Based on observations made with the Nordic Optical Telescope (NOT), owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University at the Observatorio del Roque de los Muchachos, La Palma, Spain, of the Instituto de Astrofisica de Canarias. The data presented here were obtained in part with ALFOSC, which is provided by the Instituto de Astrofisica de Andalucia (IAA) under a joint agreement with the University of Copenhagen and NOT. Some of the data were obtained during CAT service observation Spanish time. ALFOSC polarimetry imaging data of SN2020ank and calibration data were retrieved from the NOT public archive.
The Liverpool Telescope is operated on the island of La Palma by Liverpool John Moores University in the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofisica de Canarias with financial support from the UK Science and Technology Facilities Council.
This work has made use of data from the Asteroid Terrestrial- impact Last Alert System (ATLAS) project. The Asteroid Terrestrial- impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile.
Lasair is supported by the UKRI Science and Technology Facilities Council and is a collaboration between the University of Edinburgh (grant ST/N002512/1) and Queen’s University Belfast (grant ST/N002520/1) within the LSST:UK Science Consortium.
SNID is Copyright (C) 1999-2007 Stéphane Blondin and John L. Tonry, and is available under the GNU General Public License.
This research made use of the Transient Name Server (TNS) which is the official IAU mechanism for reporting new astronomical transients such as supernova candidates, As of January 1, 2016.
This research made use of Legacy Survey Data Release 10 data. The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID #2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID #2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID #2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. The Legacy Surveys project is honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation.
NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation.
This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas AM University, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo a Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Cientifico e Tecnologico and the Ministerio da Ciencia, Tecnologia e Inovacao, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energeticas, Medioambientales y Tecnologicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenossische Technische Hochschule (ETH) Zurich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciencies de l’Espai (IEEC/CSIC), the Institut de Fisica d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig Maximilians Universitat Munchen and the associated Excellence Cluster Universe, the University of Michigan, NSF’s NOIRLab, the University of Nottingham, the Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas AM University.
The Legacy Surveys imaging of the DESI footprint is supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH1123, by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract; and by the U.S. National Science Foundation, Division of Astronomical Sciences under Contract No. AST-0950945 to NOAO.
Data Availability
For science reproducibility purposes, the photometry data displayed in section LABEL:photometry_table will be available online. The spectra presented in this work are available via WISEReP. Other data can be obtained from the corresponding author upon a reasonable request.
@@ Appendix A Host Photometry Table
The measurements in the different bands of the host galaxy of SN 2021lwz discussed in Section2.1, are given in Table 4.
@@ Appendix D Linear Polarimetry material
The log of the polarimetry obtained on SN 2021lwz using ALFOSC on the NOT (see Section 2.4) is presented in Table,7. Linear polarimetry of SN 2021lwz obatined with ALFOSC in one of the V-band is shown in Figure .
@@ Appendix E Fits obtained with SNID
Appendix F Spectrum modeling with SYN++
The best-fit global parameter values of the SYN++ modeling of SN 2021lwz (see Section 4.1 )are given in Table,1. The best-fit local SYN++ parameters of the modeled spectra of SN 2021lwz are given in this appendix in Table,9.
| Ions | He i | C ii | O i | Na i | Mg ii | Si ii | P iii | Ca ii | Ti ii | Fe ii |
| -3.4 days phase | ||||||||||
| -1.0 | -1.0 | 0.9 | -0.5 | -0.4 | -0.3 | -0.3 | -1.6 | -0.5 | ||
| 15.0 | 15.0 | 15.0 | 15.0 | 15.0 | 15.0 | 15.0 | 15.0 | 15.0 | ||
| 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | ||
| aux | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | |
| 20.0 | 20.0 | 5.0 | 20.0 | 20.0 | 20.0 | 20.0 | 7.0 | 12.0 | ||
| +13.2 days and +29.1 days phase | ||||||||||
| -0.3 | -0.8 | -0.3 | 0.3 | 0.0 | 0.0 | 0.1 | ||||
| 7.5 | 7.5 | 7.5 | 7.5 | 7.5 | 7.5 | 7.5 | ||||
| 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | ||||
| aux | 10.0 | 5.0 | 3.0 | 1.0 | 2.0 | 5.0 | 3.0 | |||
| 8.0 | 20.0 | 8.0 | 8.0 | 20.0 | 15.0 | 8.0 | ||||
Appendix G Magnetar-driven model of SN 2021lwz
We fit the mulitband light curve of SN 2021lwz shown in Figure 12 using the software, model, and priors discussed in Section 4.3. The posterior is shown in Figure 21.
Appendix H Magnetar-driven model of iPTF16asu
We fit the mulitband light curve of iPTF16asu using the same software, model, and priors as for SN 2021lwz (see Section 4.3 for details). The light curve fit is shown in Figure 20 and posterior is shown in Figure 21.
Appendix I USSNe Table
The list of USSNe objects displayed in and their associated properties are given in Table 10.

### v2

@@ 5.7 Comparisons with fast evolving luminous Type Ic-BL SN 2018gep
All in all, all these comparisons suggest that SN 2021lwz has characteristics of, and is probably closely related to, many of these classes, and is therefore most likely some kind of transitional event.
@@ 5.8 Prospects and final remarks
An interesting test of the magnetar scenario for luminous, fast evolving objects is non-thermal emission. These objects could emit luminous X-ray (Kotera et al. 2013; Metzger et al. 2014; Murase et al. 2015) and radio emission (Murase et al. 2015; Omand et al. 2018; Eftekhari et al. 2021; Murase et al. 2021; Omand et al. 2025a), similar to PTF 10hgi (Eftekhari et al. 2019; Law et al. 2019; Mondal et al. 2020) or SN 2012au (Stroh et al. 2021) after a few years, but on timescales of a few weeks to a year. There could also be bright infrared emission on short timescales due to dust heated by the PWN (Omand et al. 2019), although this emission would likely fade quickly as the dust cools. The dominant non-thermal component can be predicted, in principle, from the nebular spectrum of the supernova, since line emission is sensitive to the underlying PWN spectrum (Omand and Jerkstrand 2023); however, these models are not currently capable of reproducing nebular spectra on timescales of less than a year.
The multi-dimensional structure of these magnetar-driven supernovae could show strong clumping due to Rayleigh-Taylor instabilities between the nebula and ejecta (Blondin and Chevalier 2017; Suzuki and Maeda 2017). When , such as in SN 2021lwz, the shock can break through the ejecta (Blondin and Chevalier 2017), leading to a decrease in the ejecta acceleration efficiency. This process, known as blowout, has been inferred to happen in the Crab Nebula (Omand et al. 2025b). Clumpy ejecta can cause polarization if the clump size is large enough (Tanaka et al. 2017), but it is currently unknown how to determine the clump size distribution from the inferred engine properties of the supernova.
It has been proposed by Zou and Cheng (2018) that Type I SLSNe and Type Ic-BL could be produced by jets. If so, they would both show strong polarization from early times. We point out that for SN 2021lwz our non-detections contribute to rule out this scenario. Additionally, if the energy was injected into the ejecta early, most of it would accelerate the ejecta instead of contributing to the luminosity (Suzuki and Maeda 2021).
@@ 6 Conclusions
An extensive analysis and modelling of the photometry, spectroscopy and polarimetry data of the SN, followed by comparisons with SNe of well know types and with rarer transients, consistently leads to the following results:
@@
Spectroscopically, SN 2021lwz looks like a Type Ic-like SN before maximum light. After maximum light some velocity lines may be a bit larger but the post-peak spectrum looks most like a SLSN and the continuum slope (see Figure 3) is much more like a SLSN that a Ic-BL, which is usually much redder from line blanketing.
@@
The light curve and multi-band light curve modelling of SN 2021lwz suggest it shares similarities with iPTF 16asu, and with SN 2018gep. The three sources show continuum blue dominated spectra before maximum light, even though SN 2021lwz shows spectral features pointing to a Type Ic-like SNe, while iPTF 16asu and SN 2018gep don’t. A possible explanation is that iPTF 16asu and SN 2018gep are more energetic explosions quickly rising to higher photosphere temperatures than is SN 2021lwz.
In conclusion SN 2021lwz is an uncommon transient showing many similarities with several classes of transients, and with rare transients. It may be an interesting example pointing on how differences in ejecta mass and engine parameters could produce a wide range of engine-driven SESNe.
@@
Acknowledgements.
The authors would like to thank the anonymous referee for her/his insightfull and constructive comments, Sheng Yang, Anna Ho and Ósmar Rodríguez for helpful discussions and inputs. We thank Rosa Clavero and David Morate for their kind support on the NOT. We thank George Khorenzhev (Space Research Institute, Russian Academy of Sciences) for contributing to some observations. F.P. and C.G.C. acknowledge support from the Spanish Ministerio de Ciencia, Innovación y Universidades (MICINN) under grant numbers PID2022-141915NB-C21. R.K.T. has been supported by the NKFIH/OTKA FK-134432 grant of the National Research, Development and Innovation Office of Hungary (NKFIH). C. M. B. O. acknowledges support from the Royal Society (grant Nos. DHF-R1-221175 and DHF-ERE-221005). N. Sarin acknowledges support from the Kavli foundation. T.-W.C. acknowledges the Yushan Fellow Program by the Ministry of Education, Taiwan for the financial support (MOE-111-YSFMS-0008-001-P1). T.K. acknowledges support from the Research Council of Finland project 360274. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute of Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, IN2P3, University of Warwick, Ruhr University Bochum, Cornell University, and Northwestern University. Operations are conducted by COO, IPAC, and UW. Zwicky Transient Facility access was supported by Northwestern University and the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA). The ZTF forced-photometry service was funded under the Heising-Simons Foundation grant No. 12540303 (PI: Graham). SED Machine is based upon work supported by the National Science Foundation under Grant No. 1106171. The Gordon and Betty Moore Foundation, through both the Data-Driven Investigator Program and a dedicated grant, provided critical funding for SkyPortal. Based on observations made with the Nordic Optical Telescope (NOT), owned in collaboration by the University of Turku and Aarhus University, and operated jointly by Aarhus University, the University of Turku and the University of Oslo, representing Denmark, Finland and Norway, the University of Iceland and Stockholm University at the Observatorio del Roque de los Muchachos, La Palma, Spain, of the Instituto de Astrofísica de Canarias. The data presented here were obtained in part with ALFOSC, which is provided by the Instituto de Astrofísica de Andalucia (IAA) under a joint agreement with the University of Copenhagen and NOT. Some of the data were obtained during CAT service observation Spanish time. ALFOSC polarimetry imaging data of SN2020ank and calibration data were retrieved from the NOT public archive. The Liverpool Telescope is operated on the island of La Palma by Liverpool John Moores University in the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias with financial support from the UK Science and Technology Facilities Council. This work has made use of data from the Asteroid Terrestrial- impact Last Alert System (ATLAS) project. The Asteroid Terrestrial- impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; by-products of the NEO search include images and catalogues from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. Lasair is supported by the UKRI Science and Technology Facilities Council and is a collaboration between the University of Edinburgh (grant ST/N002512/1) and Queen’s University Belfast (grant ST/N002520/1) within the LSST:UK Science Consortium. SNID is Copyright (C) 1999-2007 Stéphane Blondin and John L. Tonry, and is available under the GNU General Public License. This research made use of the Transient Name Server (TNS) which is the official IAU mechanism for reporting new astronomical transients such as supernova candidates, As of January 1, 2016. This research made use of Legacy Survey Data Release 10 data. The Legacy Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS; Proposal ID #2014B-0404; PIs: David Schlegel and Arjun Dey), the Beijing-Arizona Sky Survey (BASS; NOAO Prop. ID #2015A-0801; PIs: Zhou Xu and Xiaohui Fan), and the Mayall z-band Legacy Survey (MzLS; Prop. ID #2016A-0453; PI: Arjun Dey). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. The Legacy Surveys project is honored to be permitted to conduct astronomical research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas AM University, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo, Financiadora de Estudos e Projetos, Fundacao Carlos Chagas Filho de Amparo a Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Cientifico e Tecnologico and the Ministerio da Ciencia, Tecnologia e Inovacao, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energeticas, Medioambientales y Tecnologicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenossische Technische Hochschule (ETH) Zurich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciencies de l’Espai (IEEC/CSIC), the Institut de Fisica d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig Maximilians Universitat Munchen and the associated Excellence Cluster Universe, the University of Michigan, NSF’s NOIRLab, the University of Nottingham, the Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas AM University. The Legacy Surveys imaging of the DESI footprint is supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH1123, by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract; and by the U.S. National Science Foundation, Division of Astronomical Sciences under Contract No. AST-0950945 to NOAO.Data Availability
For science reproducibility purposes, the photometry and the spectra presented in this work are available via WISEReP191919https://www.wiserep.org/object/18321. (Yaron and Gal-Yam 2012).
@@ Appendix A Host Photometry Table
The measurements in the different bands of the host galaxy of SN 2021lwz discussed in Section2.1, are given in Table4. The magnitudes are not corrected for extinction.
@@ Appendix D Linear Polarimetry material
Linear polarimetry of SN 2021lwz obtained with ALFOSC in one of the V-band is shown in Figure 18. The log of the polarimetry obtained on SN 2021lwz using ALFOSC on the NOT (see Section 2.4) is presented in Table,7.
@@ Appendix E Fits obtained with SNID
Appendix F Spectrum modelling with SYN++
The best-fit global parameter values of the SYN++ modelling of SN 2021lwz (see Section 4.1 )are given in Table,1. The best-fit local SYN++ parameters of the modelled spectra of SN 2021lwz are given in this appendix in Table,9.
| Ions | He i | C ii | O i | Na i | Mg ii | Si ii | Ca ii | Ti ii | Fe ii |
| -3.4 days phase | |||||||||
| -1.0 | -1.0 | 0.9 | -0.5 | -0.4 | -0.3 | -1.6 | -0.5 | ||
| 15.0 | 15.0 | 15.0 | 15.0 | 15.0 | 15.0 | 15.0 | 15.0 | ||
| 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | ||
| aux | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | |
| 20.0 | 20.0 | 5.0 | 20.0 | 20.0 | 20.0 | 7.0 | 12.0 | ||
| +13.2 days and +29.1 days phase | |||||||||
| -0.3 | -0.8 | -0.3 | 0.3 | 0.0 | 0.0 | 0.1 | |||
| 7.5 | 7.5 | 7.5 | 7.5 | 7.5 | 7.5 | 7.5 | |||
| 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | 50.0 | |||
| aux | 10.0 | 5.0 | 3.0 | 1.0 | 2.0 | 5.0 | 3.0 | ||
| 8.0 | 20.0 | 8.0 | 8.0 | 20.0 | 15.0 | 8.0 | |||
Appendix G Bolometric Light curve Arnett modelling
In this section we give the details about the bolometric light curve modelling using the Arnett model (see Arnett 1982; Kasen and Bildsten 2010; Woosley 2010; West et al. 2023) which main conclusion (the light curve can’t be fit with ) is given in Section 4.2.
By re-parameterising the 56Ni mass as a fraction of the ejecta mass we then force the fit to be physical. However, as shown by the dashed red line in Figure 20, the fit converges poorly.
Appendix H Magnetar-driven model of SN 2021lwz
We fit the mulit-band light curve of SN 2021lwz shown in Figure 11 using the software, model, and priors discussed in Section 4.3. The posterior is shown in Figure 23.
Appendix I Magnetar-driven model of iPTF16asu
We fit the mulit-band light curve of iPTF16asu using the same software, model, and priors as for SN 2021lwz (see Section 4.3 for details). The light curve fit is shown in Figure 22 and posterior is shown in Figure 23.
Appendix J USSNe Table
The list of USSNe objects displayed in Section 4.4 and their associated properties are given in Table 10.